the inequality x+c>-2.55 has the solution x>4.85. What is the value of c? How do you know?


1
Expert's answer
2021-01-05T17:53:34-0500

"x+c>-2.55"

Subtracting "c" from both sides of the inequality, we get:

"x+c-c>-2.55-c"

"x>-2.55-c"

Since "x>4.85" is the solution, we conclude that "-2.55-c=4.85" .

Add "c" to both sides of the equality:

"-2.55-c+c=4.85+c"

"-2.55=4.85+c"

Subtract "4.85" from both sides of the equality:

"-2.55-4.85=4.85+c-4.85"

"-7.4=c"


Answer: "c=-7.4" .